# Fake News in Sheep’s Clothing: Robust Fake News Detection Against LLM-Empowered Style Attacks

### v1

@@ 5. Proposed Approach
Building upon our empirical findings regarding the style-related vulnerability of fake news detectors, and the limitations of LLMs as zero-shot detectors, we present SheepDog, a style-agnostic detector that reliably assesses news veracity through the utilization of LLMs. As illustrated in Figure 2, the robustness of SheepDog is obtained from two core objectives in a multi-task learning paradigm: (1) style-agnostic training, which flexibly adapts a pre-trained LM to the fake news classification objective, and ensures consistent veracity predictions across a diverse array of news reframings; and (2) content-focused veracity attributions, which extracts veracity-related insights from LLMs to inform model predictions.
To enhance usability and composability, we design our method to be simple and modular, allowing it to be used with any language model backbone.
@@ 5.1. LLM-Empowered News Reframing
In light of our Observation 1, it becomes evident that text-based fake news detectors, when exclusively trained on style-consistent news articles, may exhibit limited real-world adaptability. To overcome this limitation, our key idea is to diversify the training data through a process we refer to as reframing, which presents each news articles in various styles.
Our reframing strategy is driven by two sub-goals: (1) ensuring coherence in generated texts, and (2) covering a diverse range of styles. LLMs inherently exhibit both desirable properties. Hence, for each training article, we generate a series of prompts, each consisting of the news article and a style-oriented reframing instruction. The general prompt format is represented as follows, with a detailed example provided in Table 9:
{tcolorbox}
[colback=black!2!white,colframe=white!50!black,boxrule=0.5mm] Rewrite the following article in a / an [specified] tone: [news article] To generate news expressions that simulate both reliable and unreliable sources, we establish a set of four general style-oriented adjectives for the prompt: ”objective and professional” and ”neutral” to emulate reliable sources, and ”emotionally triggering” and ”sensational” for unreliable ones.
During the training phase, for labeled news article $p\in\mathcal{P}_{L}$, we randomly select one reliable-style reframing prompt and one unreliable-style reframing prompt to generate diverse expressions. Through querying the LLM, we obtain two corresponding reframings: one reliable-style reframing denoted as $p_{R}$, and one unreliable-style reframing denoted as $p_{F}$.
@@ 5.2. Style-Agnostic Training
A robust fake news detector should possess the capability to discern the veracity of news articles based on their content, without being influenced by their stylistic attributes. To this end, we introduce a style alignment objective that ensures close alignment among the veracity predictions of news article $p$, its reliable-style reframing $p_{R}$, and its unreliable-style reframing $p_{F}$. This objective is derived as follows.
@@
Subsequently, we apply a Multi-Layer Perceptron (MLP) to these representations, and obtain veracity predictions $\tilde{\mathbf{y}},\tilde{\mathbf{y}}_{R},\tilde{\mathbf{y}}_{F}\in\mathbb{R}^{2}$ for $p$, $p_{R}$ and $p_{F}$:
@@
Despite differences in style, the fundamental news content remains consistent across the original news article $p$ and its reframings $p_{R}$ and $p_{F}$. Ideally, $p_{R}$ and $p_{F}$ should yield the same predicted label as $p$. To facilitate this, we formulate the following style alignment loss defined as follows:
@@
While aligning the predictions of $p_{R}$ and $p_{F}$ with $p$, it is crucial to ensure accurate predictions for $p$. Therefore, we also incorporate a fake news detection loss:
@@ 5.3. Content-Focused Veracity Attributions
In addition to tuning the LM backbone with the style alignment objective for robust predictions, to capture the veracity-related signals directly linked to the news content, we further propose to integrate external knowledge and reasoning to inform veracity predictions. General-purpose LLMs, known for their remarkable zero-shot reasoning capabilities (Huang and Chang, 2023), provides a promising solution.
Specifically, we generate content-focused veracity attributions by prompting an LLM. Our prompt consists of the news article text and a predefined set $\mathcal{C}$ of content-oriented fact-checking rationales (e.g., “lack of credible sources” and “false or misleading information”; we present the detailed prompt template and an example in Appendix B.3 and Table 10). Through this prompt, we instruct the LLM to provide textual attributions indicative of news veracity (i.e., identifying characteristics associated with fake news). This process efficiently extracts the LLM’s reasoning capabilities and general knowledge, presenting it in a text format. The resulting textual attributions, in turn, enriches our model with additional veracity-informed knowledge. The general prompt format is shown as follows:
{tcolorbox}
[colback=black!2!white,colframe=white!50!black,boxrule=0.5mm] Article: [news article]
Question: [given a list of fact-checking rationales, ask the LLM to identify content-related characteristics associated with fake news] Upon querying the LLM, we obtain a list of textual veracity attributions based on the input article. These attributions are then transformed into $|\mathcal{C}|$-dimensional pseudo-labels, where the attribution to each rationale in $\mathcal{C}$ is represented by a distinct binary label. For a given news article $p\in\mathcal{P}_{L}$, this process yields pseudo-labels $\mathbf{s}\in\mathbb{R}^{|\mathcal{C}|}$, which provides supplementary veracity-related information. Similarly, for reframings $p_{R}$ and $p_{F}$, we obtain pseudo-labels $\mathbf{s}_{R},\mathbf{s}_{F}\in\mathbb{R}^{|\mathcal{C}|}$, respectively. Notably, as $\mathcal{C}$ is focused on indicators of fake news, the pseudo-labels for real news and its reframings are uniformly set to all-zeros.
To incorporate the veracity-informed knowledge from these content-focused veracity attributions into our detector, we introduce a multi-label attribution prediction objective. As shown in Eq. 1, we learn news representations $\mathbf{h},\mathbf{h}_{R},\mathbf{h}_{F}\in\mathbb{R}^{d}$ for news article $p$ and its reframings $p_{R}$ and $p_{F}$, respectively, using a pre-trained LM $\mathcal{M}$. Then, the attribution-level prediction scores $\tilde{\mathbf{s}},\tilde{\mathbf{s}}_{R},\tilde{\mathbf{s}}_{F}\in\mathbb{R}^{|\mathcal{C}|}$ are computed through another MLP:
@@ 5.4. Final Objective Function of SheepDog
Our proposed SheepDog detector is an end-to-end framework where we simultaneously learn the style-agnostic news veracity predictor and the content-focused attribution predictor.
@@ A B C D A B C D A B C D
Ours SheepDog $*{}^{*}$ $*{}^{*}$ $*{}^{*}$ 81.24 $*{}^{*}$ $*{}^{*}$ $*{}^{*}$ $*{}^{*}$ $*{}^{*}$ $*{}^{*}$ $*{}^{*}$ $*{}^{*}$
@@ 6. Experiments
  • Effectiveness on Unperturbed Articles (Section 6.3): How proficiently can SheepDog identify fake news within the original unperturbed test articles?
  • Adaptability to Different LM Backbones (Section 6.4): How does SheepDog’s performance vary with different representative LM backbones?
  • Ablation Study (Section 6.5): What are the respective impacts of style-agnostic training and content-focused attributions on SheepDog’s overall performance?
  • Stability Across Reframing Prompts (Section 6.6): Does SheepDog yield consistent improvements across diverse settings of news reframing prompts?
  • Case Study (Section 6.7): How can we interpret SheepDog’s rationale for fake news detection through its predictions on content-focused attributions?
@@ 6.1.1. Datasets
We evaluate our approach on three widely-used real-world benchmark datasets: the FakeNewsNet public benchmark (Shu et al., 2020), which consists of the PolitiFact and GossipCop datasets, and the Labeled Unreliable News (LUN) dataset (Rashkin et al., 2017). PolitiFact and LUN center on political discourse, while GossipCop focuses on celebrity gossip. PolitiFact and LUN center on political discourse, while GossipCop focuses on celebrity gossip.
The dataset description and statistics are delegated to Appendix A and Table 8. We randomly select $80$% of the articles as the training data, and the other $20$% as test data.
@@ 6.1.2. Baselines
LLMs (G3) generate textual veracity predictions in a zero-shot manner. We select three representative LLMs as baselines: GPT3.5 (OpenAI, 2022) (version name: gpt-3.5-turbo), InstructGPT (Ouyang et al., 2022) (version name: gpt-3.5-turbo-instruct), and Llama2-13B (Touvron et al., 2023) (version name: Llama-2-13b-chat-hf), with details delegated to Appendix B.1.
@@ 6.1.3. Implementation Details
We implement SheepDog and its variants based on PyTorch 1.10.0 with CUDA 11.1. We utilize pretrained RoBERTa-base weights from HuggingFace Transformers 4.13.0 (Wolf et al., 2020). The LM backbone for SheepDog was configured with a maximum sequence length of 512, a batch size of 4, and a learning rate of $2\times 10^{-5}$. We prompt GPT3.5 (version name: gpt-3.5-turbo) to generate news reframings, content-focused veracity attributions, and adversarial test articles. Detailed prompting descriptions and examples are delegated to Appendix B.2 and B.3. For SheepDog’s attribution prediction and veracity prediction, we employ two MLPs, each with a single layer (we also implement a variant with 2-layer MLPs in Section 6.5). Our model is fine-tuned for $5$ epochs. For the implementation of baseline methods, we adhere to the architectures and hyperparameters recommended by their respective authors.
@@ 6.2. Robustness Against Style Attacks
We establish a series of LLM-empowered style attacks to assess SheepDog’s robustness. Following our prompt template formulated in Section 4.1, in the place of [publisher name], we select “National Enquirer” and “The Sun” to camouflage real news, and “CNN” and “The New York Times” for fake news, according to publisher popularity. This yields $2\times 2=4$ distinct adversarial test sets, labeled as A through D in Table 3. Note that we report the results of SheepDog on adversarial set A in all other subsections.
[publisher name] CNN The New York Times
National Enquirer A B
The Sun C D
@@
6.3. Effectiveness on Unperturbed Articles
A desirable fake news detector should achieve both robustness and effectiveness. Our empirical results, detailed in Table 4, clearly demonstrate that SheepDog excels in this regard. When tested on the original, unaltered articles, SheepDog consistently matches (on PolitiFact and GossipCop) or surpasses (on LUN) the performance of the most competitive baseline, in terms of accuracy and F1 score. This showcases the framework’s proficiency in accurately discerning unreliable news from reliable ones, and implies SheepDog’s practical utility in real-world scenarios.
6.4. Adaptability to Different LM Backbones
To assess the flexibility of SheepDog, we evaluate the performance of SheepDog combined with three representative LM backbones: RoBERTa, BERT and DeBERTa. As demonstrated in Table 5, SheepDog substantially enhances the performance of the corresponding LM backbones. This adaptability highlights SheepDog’s robustness from style-agnostic training and content-focused attributions, showcasing its potential for flexible deployment in diverse contexts where different LMs may be preferred or more readily available. This finding further reinforces the benefits of SheepDog’s design with a LM backbone, which flexibly adapts it to task-specific needs.
@@ 6.4. Adaptability to Different LM Backbones
SheepDog-RoBERTa $*{}^{*}$ $*{}^{*}$ $*{}^{*}$
BERT 72.31 68.98 56.35
SheepDog-BERT $*{}^{*}$ $*{}^{*}$ $*{}^{*}$
@@
SheepDog-DeBERTa $*{}^{*}$ $*{}^{*}$ $*{}^{*}$
@@ 6.5. Ablation Study
To gain deeper insights into the functioning of SheepDog and the role of its different components, we conduct an ablation study to evaluate the respective contributions of LLM-empowered news reframings and content-focused veracity attributions. We compare SheepDog with the following three model variants:
@@
  • SheepDog-A, which omits content-focused veracity attributions and the attribution prediction component.
Table 6 shows the performance of different SheepDog variants on adversarial test sets C and D. The results suggest that: (1) SheepDog-R without news reframings only yields slight improvements compared with a fine-tuned RoBERTa model, which suggests the pivotal role of diverse reframings in SheepDog’s robustness. (2) Removing content-focused veracity attributions (SheepDog-A) results in inferior performance than SheepDog, which implies the effectiveness of extracting content-related knowledge from LLMs. (3) SheepDog, utilizing one single layer for veracity prediction and attribute predictions, slightly outperforms the variant employing 2-layer MLPs. This suggests that the expressiveness of LMs, harnessed through our proposed objectives, yields article representations that contains rich indicators for both attributions and veracity.
@@ 6.6. Stability Across Reframing Prompts
As detailed in Section 5.1, we utilize an LLM to generate two types of reframings for a given training article $p$: a reliable-style reframing denoted as $p_{R}$, and an unreliable-style reframing denoted as $p_{F}$. To assess SheepDog’s stability across different reframing prompts, we systematically examine its performance using four diverse combinations of prompts, denoted as R1 through R4.
@@
{tcolorbox}
[colback=black!2!white,colframe=white!50!black,boxrule=0.5mm] Rewrite the following article in a / an [specified] tone: [$p$]
For R1 through R4, the specified tones are defined as:
  • R1: $p_{R}$ “objective and professional”; $p_{F}$ “emotionally triggering”.
  • R2: $p_{R}$ “objective and professional”; $p_{F}$ “sensational”.
  • R3: $p_{R}$ “neutral”; $p_{F}$ “emotionally triggering”.
  • R4: $p_{R}$ “neutral”; $p_{F}$ “sensational”.
As shown in Table 7, SheepDog consistently demonstrates stable and significant improvements over the most competitive baseline. This validates the generalizability of our approach, suggesting its potential in effectively combating deceptive online information in the ever-evolving digital landscape. Notably, our SheepDog approach conducts sampling between two reliable-style reframings and two unreliable-style reframings, leading to better versatility.
Table 7. Across different sets of news reframing prompts, SheepDog demonstrates stable and significant improvements over the most competitive baseline (in F1 %). (“Baseline”: performance of the most competitive baseline.
@@
Figure 3. SheepDog yields robust predictions against LLM-camouflaged fake news. Additionally, the top predicted veracity attribution points toward the false and misleading information in the news content, providing interpretability.

### v2

@@ 5. Proposed Approach
Building upon our empirical findings on the style-related vulnerability of text-based fake news detectors, we introduce SheepDog, a style-agnostic detector that reliably assesses news veracity. As overviewed in Figure 2, SheepDog obtains style robustness from two core objectives within a multi-task learning paradigm: (1) style-agnostic training, which flexibly adapts an LM to the fake news classification task, while ensuring consistent veracity predictions across a diverse array of LLM-empowered news reframings; and (2) content-focused veracity attribution prediction, which leverages veracity-related insights from LLMs to inform model predictions. To enhance usability and composability, we design our method to be simple and modular, allowing it to be integrated with any LM and LLM backbone.
@@ 5.1. LLM-Empowered News Reframing
As suggested by our Observation 1, text-based fake news detectors fitted to style-consistent real and fake articles exhibit limited adaptability against stylistic variations. To overcome this limitation, our key idea is to inject style diversity into the training stage through a process we term reframing, where each news article is presented in various styles.
SheepDog’s reframing strategy is driven by two sub-goals: (1) encompassing a wide range of styles and (2) maintaining the integrity of the original news content. LLMs, capable of following complex real-world instructions (He et al., 2024), inherently meet both criteria. This is further validated by our analysis in Appendix C, where LLMs generally prove effective in transforming the tone of news articles while preserving content consistency. Hence, for each training article, we generate a series of prompts, each comprising the news article and a style-oriented reframing instruction. The general structure of the prompt is as follows, with a detailed example presented in Table 13:
Rewrite the following article in a / an [specified] tone: [news article]
To generate news expressions that simulate both reliable and unreliable sources, we establish a set of four general style-oriented adjectives for the prompt: “objective and professional” and “neutral” to emulate reliable sources, and “emotionally triggering” and “sensational” for unreliable sources. During the training stage, for labeled news article $p\in\mathcal{P}_{L}$, we randomly select one reliable-style reframing prompt and one unreliable-style reframing prompt to generate diverse expressions. Through querying the LLM, we obtain two corresponding reframings: one reliable-style reframing denoted as $p_{R}$, and one unreliable-style reframing denoted as $p_{F}$.
@@ 5.2. Style-Agnostic Training
A style-robust fake news detector must be capable of discerning the veracity of news articles based on their content, without being influenced by stylistic features. To this end, we introduce a style alignment objective that ensures close alignment among the veracity predictions of news article $p$, its reliable-style reframing $p_{R}$, and its unreliable-style reframing $p_{F}$. This objective is derived as follows.
@@
Subsequently, we apply a Multi-Layer Perceptron (MLP) to $p$, $p_{R}$ and $p_{F}$ to obtain corresponding veracity predictions $\tilde{\mathbf{y}},\tilde{\mathbf{y}}_{R},\tilde{\mathbf{y}}_{F}\in\mathbb{R}^{2}$:
@@
Despite differences in style, the fundamental news content remains consistent across the original news article $p$ and its reframings $p_{R}$ and $p_{F}$. Ideally, $p_{R}$ and $p_{F}$ should yield the same veracity prediction as $p$. To this end, we formulate the following style alignment loss defined as:
@@
While aligning the predictions of $p_{R}$ and $p_{F}$ with $p$, it is crucial to ensure accurate veracity prediction for $p$. Therefore, we also incorporate a fake news detection loss:
@@ 5.3. Content-Focused Veracity Attributions
In addition to tuning the LM backbone with the style alignment objective, which discounts style-related features and encourages style-robust predictions, we further propose to integrate auxiliary veracity-related knowledge and reasoning to inform veracity predictions. To achieve this goal, leveraging the impressive zero-shot reasoning capabilities of general-purpose LLMs (Asai et al., 2024; Menon and Vondrick, 2023) serves as a promising solution.
Specifically, we elicit content-focused veracity attributions from an LLM, which provides explanatory outputs on why each fake news article in the training set is flagged as fake. Our prompt consists of a fake news article and a predefined set $\mathcal{C}$ of content-oriented rationales for debunking fake news (e.g., “lack of credible sources” and “false or misleading information”; detailed rationales are described in Appendix B.3). This prompt efficiently leverages the LLM’s reasoning capabilities and prior knowledge to identify characteristics associated with fake news, referencing the rationales in $\mathcal{C}$. The general prompt format is as follows:
Article: [fake news article] Question: [given a list of content-centric rationales for debunking fake news, ask the LLM to identify rationales fulfilled by the fake news article]
Upon querying the LLM, we obtain a list of veracity attributions based on the input article. These attributions are then converted into $|\mathcal{C}|$-dimensional pseudo-labels, where each rationale in $\mathcal{C}$ is represented by a distinct binary label. For a given news article $p\in\mathcal{P}_{L}$, this process yields pseudo-labels $\mathbf{s}\in\mathbb{R}^{|\mathcal{C}|}$, which contains supplementary veracity-related information. Similarly, for reframings $p_{R}$ and $p_{F}$, we obtain pseudo-labels $\mathbf{s}_{R},\mathbf{s}_{F}\in\mathbb{R}^{|\mathcal{C}|}$, respectively. Notably, since $\mathcal{C}$ focuses solely on fake news indicators, the pseudo-labels for real news and its reframings are uniformly set to all zeros.
To distill the veracity-informed knowledge from these attributions, we introduce a multi-label attribution prediction objective. This enriches our framework with additional content-centric guidance, and offers potential explanability for articles identified as fake during the inference stage (exemplified in Figure 3).
As shown in Eq. 1, we learn news representations $\mathbf{h},\mathbf{h}_{R},\mathbf{h}_{F}\in\mathbb{R}^{d}$ for news article $p$ and its reframings $p_{R}$ and $p_{F}$, respectively, using a pre-trained LM $\mathcal{M}$. Then, the attribution-level prediction scores $\tilde{\mathbf{s}},\tilde{\mathbf{s}}_{R},\tilde{\mathbf{s}}_{F}\in\mathbb{R}^{|\mathcal{C}|}$ are computed through another MLP:
@@ 5.4. Final Objective Function of SheepDog
SheepDog is designed as an end-to-end framework, where the style-agnostic news veracity predictor and the content-focused attribution predictor are trained simultaneously.
Table 2. Dataset statistics.
Dataset PolitiFact GossipCop LUN
# News Articles 450 7,916 7,500
# Real News 225 3,958 3,750
# Fake News 225 3,958 3,750
@@ A B C D A B C D A B C D
Ours SheepDog 80.99 79.89 82.36 81.24 74.45 74.38 75.95 75.88 85.63 86.06 87.89 88.32
[publisher name] CNN The New York Times
National Enquirer A B
The Sun C D
@@ 6. Experiments
  • Effectiveness on Unperturbed Articles (Section 6.3): How effectively can SheepDog identify fake news within the original unperturbed test articles?
  • Adaptability to Different Backbones (Section 6.4): How well does SheepDog perform when combined with different LM and LLM backbones?
  • Ablation Study (Section 6.5): What are the respective roles of style-agnostic training and content-focused attributions on SheepDog’s style robustness?
  • Stability Across Reframing Prompts (Section 6.6): Does SheepDog yield consistent improvements across diverse sets of news reframing prompts?
  • Case Study (Section 6.7): How can we interpret SheepDog’s rationale for debunking fake news through its predictions on content-focused veracity attributions?
@@ 6.1.1. Datasets
We evaluate our approach on three widely-used real-world benchmark datasets: the FakeNewsNet public benchmark (Shu et al., 2020), which consists of the PolitiFact and GossipCop datasets, and the Labeled Unreliable News (LUN) dataset (Rashkin et al., 2017). Table 2 describes the dataset statistics. PolitiFact and LUN center on political discourse, while GossipCop focuses on celebrity gossip. For the LUN dataset, which further classifies unreliable news into three sub-categories: satire, hoax, and propaganda, we conduct binary classification between reliable (real) and unreliable (fake) news, and ensure an equal number of unreliable news from each of these fine-grained categories. To better simulate real-world scenarios, we follow prior work (Wu and Hooi, 2023) and adopt temporal data splitting on PolitiFact and GossipCop, where temporal information is available. The most recent 20% real and fake news articles constitute the test set, and the remaining 80% articles posted earlier serve as the training set. We adopt random 80/20 training / test splits on LUN.
@@ 6.1.2. Baselines
LLMs (G3) conduct zero-shot veracity prediction. We select three representative baseline LLMs: GPT-3.5 (OpenAI, 2022), InstructGPT (Ouyang et al., 2022), and LLaMA2-13B (Touvron et al., 2023), detailed in Appendix B.1.
@@ 6.1.3. Implementation Details
We implement SheepDog and its variants based on PyTorch 1.10.0 with CUDA 11.1. We utilize pretrained RoBERTa-base weights from HuggingFace Transformers 4.13.0 (Wolf et al., 2020). The LM backbone for SheepDog was configured with a maximum sequence length of 512, a batch size of 4, and a learning rate of $2\times 10^{-5}$. We prompt GPT-3.5 to generate news reframings, content-focused veracity attributions, and adversarial test articles. (detailed prompting descriptions and examples are presented in Appendix B.2 and B.3). For SheepDog’s attribution prediction and veracity prediction, we employ two MLPs, each with a single layer (we also implement a variant with 2-layer MLPs in Section 6.5). Our model is fine-tuned for $5$ epochs. For the implementation of baseline methods, we adhere to the architectures and hyperparameters recommended by their respective authors.
@@ 6.2. Robustness Against Style Attacks
We establish a series of LLM-empowered style attacks to assess SheepDog’s robustness. Following our prompt template formulated in Section 4.1, in the place of [publisher name], we select “National Enquirer” and “The Sun” to camouflage real news, and “CNN” and “The New York Times” for fake news, according to publisher popularity. This yields $2\times 2=4$ distinct adversarial test sets, labeled as A through D in Table 4. Note that we report the results of SheepDog on adversarial set A in all other subsections, unless otherwise specified.
@@
SheepDog-RoBERTa 80.99 74.45 85.63
BERT 72.31 68.98 53.97
SheepDog-BERT 81.37 73.54 80.36
@@
SheepDog-DeBERTa 81.10 73.89 82.58
6.3. Effectiveness on Unperturbed Articles
A desirable fake news detector should achieve style robustness under adversarial settings without compromising its effectiveness under the unperturbed setting. Our empirical results, presented in Table 5, demonstrate that SheepDog excels in this regard. When tested on the original, unaltered articles, SheepDog consistently matches (on PolitiFact and GossipCop) or surpasses (on LUN) the performance of the most competitive baseline, in terms of both accuracy and F1 score. Similar to our observation (4) in Section 6.2, the significant performance gains on LUN might also stem from dataset-specific stylistic features (detailed in Section 6.8).
6.4. Adaptability to LM / LLM Backbones
To assess the flexibility of SheepDog, we evaluate the performance of SheepDog combined with three representative LMs: RoBERTa, BERT and DeBERTa. We also evaluate RoBERTa-based SheepDog combined with two representative LLMs: the closed-source GPT-3.5 and the open-source LLaMA2-13B.
As demonstrated in Table 6, SheepDog (1) substantially enhances the performance of each respective LM backbone. Additionally, as shown in Table 7, SheepDog (2) also achieves superior style robustness when utilizing both closed-source and open-source LLMs. This adaptability highlights SheepDog’s style robustness from style-invariant training and content-focused attribution prediction, implying its practical utility in real-world scenarios where different LMs and LLMs may be preferred or more readily available.
[…]
Method PolitiFact GossipCop LUN
SheepDog 80.99 74.45 85.63
SheepDog-LLaMA2 80.82 74.04 81.87
RoBERTa 76.17 71.00 52.47
@@ 6.5. Ablation Study
To gain deeper insights into the functioning of SheepDog and the role of its different components, we compare SheepDog with the following three model variants:
@@
  • SheepDog-A, which excludes content-focused veracity attributions and the attribution prediction component.
Results in Table 8 suggest that: (1) SheepDog-R without news reframings only yields slight improvements over fine-tuned RoBERTa, which suggests the key role of diverse reframings in SheepDog’s robustness. (2) While the improvement of SheepDog over SheepDog-A may seem slight, incorporating veracity attributions assists in guiding the model to prioritize content over style. Furthermore, SheepDog’s attribution predictor equips the framework with explanatory outputs during inference stage, facilitating easier human verification in real-world scenarios (see Figure 3 and Section 6.7 for a concrete illustration of this functionality). (3) SheepDog, utilizing one single layer for veracity prediction and attribution predictions, slightly outperforms the variant employing 2-layer MLPs. This suggests that the expressiveness of LMs, harnessed through our proposed objectives, yields article representations that contain rich indicators related to both attributions and veracity.
@@ 6.6. Stability Across Reframing Prompts
As detailed in Section 5.1, we utilize an LLM to generate two types of reframings for a given training article $p$: a reliable-style reframing denoted as $p_{R}$, and an unreliable-style reframing denoted as $p_{F}$. To assess SheepDog’s stability across different reframing prompts, we examine its performance using four diverse combinations of prompts, denoted as R1 through R4.
@@
Rewrite the following article in a / an [specified] tone: [$p$]
For R1 through R4, the specified tones are defined as: ($p_{R}$ / $p_{F}$)
  • R1: “objective and professional” /“emotionally triggering”.
  • R2: “objective and professional” / “sensational”.
  • R3: “neutral” / “emotionally triggering”.
  • R4: “neutral” / “sensational”.
As shown in Table 9, SheepDog consistently demonstrates stable and significant improvements over the most competitive baseline. This validates the generalizability of our approach, suggesting its potential in effectively combating deceptive information in the ever-evolving digital landscape. Notably, our SheepDog approach conducts sampling between two reliable-style reframings and two unreliable-style reframings, leading to stronger versatility.
Table 9. Across different sets of news reframing prompts, SheepDog demonstrates stable and significant improvements over the most competitive baseline (in F1 %).
@@
Figure 3. Across the original fake news article and its LLM-camouflaged counterparts, SheepDog maintains consistency and accuracy in both its veracity prediction and the top-predicted veracity attribution for debunking fake news.